OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 0 −3 3 2

OpenStudy (anonymous):

Hint: Substitution method. Isolate y in the first equation: 4x + 2y = 12 2y = -4x + 12 y = -2x + 3 Substitute this in for the value of y in the second equation: x – y = 3 x - (2x + 3) = 3 x - 2x - 3 = 3 -x - 3= 3 -x = 6 x = 6 /-1

hartnn (hartnn):

or multiply second equation by 2 and add both the equations......

OpenStudy (anonymous):

^That is a much easier way. (: